The length of a rectangle is 7 inches greater than its width. The perimeter of the rectangle is 50 inches. Find the length and width of the rectangle.
Answer by josgarithmetic(39801)

w width
L length

L=7+w and 2w+2L=50 same as w+L=25.
Substitute.
Solve.
Substitute for result.
Solve.
Done.